Calculate the volume, in cubic centimeters, occupied by 0.022 mmole lead (density = 11.3 g/cc)
1
Expert's answer
2020-10-23T09:15:41-0400

0.022 mmole of Pb (molar mass is 207.20):

207.20 x 0.022 / 1000 = 0.0045584 g

d = m/V

V = m/d

V = 0.0045584 g / 11.3 cc = 4.0x10-4 cc (cubic cm)